OpenStudy (anonymous):

what is the solution for 2x^2+7x-2=-x^2+2x

OpenStudy (anonymous):

help

OpenStudy (jennychan12):

move all the x^2 and the x terms to one side.

OpenStudy (anonymous):

i got 1/3

OpenStudy (anonymous):

or -1/2

OpenStudy (debbieg):

\(2x^2+7x-2=-x^2+2x\) \(3x^2+5x-2=0\) \((3x-1)(x+2)=0\)